A company currently pays a dividend of $3.8 per share (D0 = $3.8). It is estimated that the company's dividend will grow at a rate of 19% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company's stock has a beta of 1.6, the risk-free rate is 7%, and the market risk premium is 3.5%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

For calculating the current market price of a share we have to find discount rate which is the cost of equity(Re).

It can be calculated using the following formula:-

Re = Rf+(Rm-Rf) beta

Where

Re= required rate of return

Rm-Rf= market risk premium

Rf= risk free rate

So, Re= 7%+(3.5)1.6= 12.6%

Now we can find the current market price of share using the following steps:-

1. Explicit forecast period (year 1-2)

2. Horizon period (after 2 years)

3. Conclusion

1. Explicit forecast period (year 1-2)

Here we have to find the present value of the the cash flow for two years. For that we can use the following formula:-

= D1/(1+Re)1 + D2/(1+Re)2

D1 and D2 means dividend for first and second years

= (3.8*1.19)/1.126 + (3.8*1.19*1.19) / (1.126*1.126)

= 4.01 + 4.244

= 8.254

2. Horizon period (after 2 years)

Value at the end of second year= D3/Re-g

= (3.8*1.19*1.19*1.05)/(0.126-0.05)

= 74.345

Present value of 74.345 in year 0= 74.345/1.1262

= 58.637

Current current price of share= 8.254+58.637= 66.891

.... Hope you are satisfied with the solution, kindly support by rating the answers.....